OpenStudy (anonymous):

what is the solution set √2-x=x

OpenStudy (kc_kennylau):

Firstly get rid of the root by squaring both sides.

OpenStudy (anonymous):

√(2-x)=x square both sides (raise each side to the second power)

OpenStudy (anonymous):

For example, if you had the following problem.... √ ( 4-b ) = b Solution: √ ( 4-b ) = b raise each side to the second power.... raising the left side to the second power: √( 4-b ) --> √( 4-b ) ² = 4 - b raising the right side to the second power: b ---> b² So, after you raise both sides to 2nd power, you will get 4 - b = b²

OpenStudy (kc_kennylau):

@tracyvsavage: student_basil has provided an example for you to follow, please respond.